### v1

# Pulse-train propagation in nonlinear Kerr media governed by higher-order dispersion

Abstract
We discover three novel classes of pulse-train waveforms in an optical Kerr nonlinear medium possessing all orders of dispersion up to the fourth order. We show that both single- and double-humped pulse-trains can be formed in the nonlinear medium. A distinguishing property is that these structures have different amplitudes, widths and wavenumbers but equal velocity which depends on the three dispersion parameters. More importantly, we find that the relation between the amplitude and duration of all the newly obtained pulse-trains is determined by the sign of a joint parameter solely. The results show that those optical waves are general, in the sense that no specified conditions on the material parameters are assumed. Considering the long-wave limit, the derived pulse-trains degenerate to soliton pulses of the quartic and dipole kinds.
pacs:
05.45.Yv, 42.65.Tg
I Introduction
The transmission of soliton pulses in nonlinear media has been the focus of continuing interest, stimulated by their important potential applications in long-distance optical communication systems and optical switching devises [1]. The formation of such wave packets in a monomode optical fiber results from the exact balance between the pulse-spreading effects of group-velocity dispersion and self-phase modulation effects due to the intensity-dependent refractive index of the fiber [2]. Experimental observation of solitons have been reported in various optical systems, including single-mode optical fibers [3], femtosecond lasers [4] and bulk optical materials [5].
In the picosecond regime, the propagation of light pulses inside a single-mode optical fiber was shown to be governed by the nonlinear Schrödinger equation (NLSE), which is commonly derived with use of the slowly varying envelope approximation [8]. This physically relevant model allows for either bright or dark solitons in the anomalous and normal dispersion regimes, respectively [9]. However, for much shorter light pulses, there is also appearance of additional higher-order effects in the optical system. For example, the third-order dispersion significantly influence the short pulses whose widths are nearly 50 femtosecond as they propagate through the nonlinear medium [10, 11]. The fourth-order dispersion is also found to be important when pulses are shorter than 10 femtosecond [10, 11]. In such a situation, extensions of the NLSE model that take into account the influence of diverse physical phenomena on short-pulse propagation and generation are therefore essential to describe the wave dynamics accurately. One notes that higher-order effects introduce several new important phenomena into the system dynamics that are absent in Kerr nonlinear medium.
A great interest has been given recently to the generation of pulse-trains in various physical systems, including birefringent optical fibers [12, 13, 14, 15, 16], negative-index materials [17], Bose-Einstein condensates [18], etc. It can be relevant to note that such kind of pulse shape is interesting in itself as another type of physically relevant nonlinear waves in optical fibers besides the soliton-type waveform. These structures exhibit many interesting properties that make them potentially in problems of signal propagation in nonlinear fibers [19]. It is intersting to note that pulse-trains can be produced experimentally. For example, the formation of cnoidal wave has been recently demonstrated in a fiber laser either in the net normal or net anomalous cavity dispersion regime [20]. Importantly, such arrays of pulses may be created by the modulational instability of the continuous-wave signal [21], or from a dual-frequency pump [22]. In optics, such nonlinear waves play a pivotal role in the analysis of the data transmission in fiber-optic telecommunications links [23].
Identification of new kinds of pulse-trains in optical systems is an issue worth considering. It is pertinent to mention that the finding of such structures may lead to novel dynamical behaviors of waves and opens the possibility for experimental investigations of various wave phenomena. In this work, we demonstrate for the first time that an optical Kerr nonlinear medium exhibiting all orders of dispersion up to the fourth order supports the existence of three novel classes of pulse-trains which take either single- or double-hump forms.
The paper is structured as follows. Section II provides the analytic framework necessary to find the traveling wave solutions of the extended NLSE model that governs the transmission of extremely short pulses throught an optical fiber medium exhibiting all orders of dispersion up to the fourth order. In Section III, we derive analytical pulse-train solutions of the model and their characteristics. We also discuss here the propagation properties of obtained pulse-trains in the fiber medium. Finally, some conclusions are given in Sec. VI.
II Model and traveling waves
We consider the evolution of an optical pulse in a highly dispersive fiber medium having a Kerr nonlinearity. The governing model for the slowly varying envelope $\psi$ of the optical pulse can be described by the extended NLSE [6, 7],
$i\frac{\partial\psi}{\partial z}=\alpha\frac{\partial\psi^{2}}{\partial\tau^{2}}+i\sigma\frac{\partial\psi^{3}}{\partial\tau^{3}}-\epsilon\frac{\partial\psi^{4}}{\partial\tau^{4}}-\gamma\left|\psi\right|^{2}\psi.$ (1)
where $\tau=t-\beta_{1}z$ is the retarded time, $z$ is the longitudinal coordinate, $\alpha=\beta_{2}/2$, $\sigma=\beta_{3}/6$, $\epsilon=\beta_{4}/24$, and $\gamma$ is the nonlinear parameter. Here the parameters $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ are the first-, second-, third-, and fourth-order dispersion coefficients, respectively.
Theoretical investigations based on Eq. (1) are mainly concentrated on the existence and propagation properties of localized pulses due to their fundamental importance in the understanding of various physical phenomena in the system. For example, Kruglov and Harvey [6] demonstrated the existence of stable solitary waves with a $\mathrm{sech}^{2}$ shape for the model. Triki and Kruglov [7] discussed the dipole soliton solution of Eq. (1) in the presence the inhomogeneities of media. Cavalcanti et al. [24] analyzed the modulational instability phenomena of Eq. (1) in the region of the minimum group-velocity dispersion. Karpman et al. [25, 26] investigated the time behavior of the amplitudes, velocities, and other parameters of radiating solitons. Shagalov [27] studied the effect of the third- and fourth-order dispersions on the modulational instability. Roy et al. [29] discussed the role of third- and fourth-order dispersions in the radiation emitted by fundamental soliton pulses in the form of dispersive waves within the framework of the dimensionless form of the model (1). Kruglov [28] obtained the periodic wave solutions that take the $\mathrm{cn}^{2}$ shape for this governing equation [28]. In what follows, we present three novel exact pulse-train solutions with interesting properties for the extended NLSE (1), which are obtained without necessarily assuming a specified condition on the fiber parameters.
In order to find some interesting analytical solutions of the model (1), we consider a solution of the form
$\psi(z,\tau)=\rho(\xi)\exp[i(\kappa z-\delta\tau+\theta)],$ (2)
where $\rho(\xi)$ is a real function, $\xi=\tau-qz$ is a traveling coordinate with $q=v^{-1}$ being the inverse velocity. Also, $\kappa$ and $\delta$ are the respective real parameters representing the wavenumber and frequency shift, while $\theta$ represents the phase of the pulse at $z=0$.
Substitution of Eq. (1) into the extended NLSE (2) leads to a set of coupled differential equations:
$(\sigma+4\epsilon\delta)\frac{d^{3}\rho}{d\xi^{3}}+(q-2\alpha\delta-3\sigma\delta^{2}-4\epsilon\delta^{3})\frac{d\rho}{d\xi}=0,$ (3)
$\epsilon\frac{d^{4}\rho}{d\xi^{4}}-(\alpha+3\sigma\delta+6\epsilon\delta^{2})\frac{d^{2}\rho}{d\xi^{2}}+\gamma\rho^{3}-(\kappa-\alpha\delta^{2}-\sigma\delta^{3}-\epsilon\delta^{4})\rho=0.$ (4)
Here, we refer the solution of the extended NLSE where the function $\psi(z,\tau)$ is given by Eq. (2) with $\rho(\xi)\neq\mathrm{constant}$ as non-plain wave or traveling wave solution.
The system of Eqs. (3) and (4) with $\epsilon\neq 0$ yields the non-plain wave solutions if and only if the next relations are satisfied:
$q=2\alpha\delta+3\sigma\delta^{2}+4\epsilon\delta^{3},~{}~{}~{}~{}\delta=-\frac{\sigma}{4\epsilon}.$ (5)
Note that Eq. (3) is satisfied for an arbitrary function $\rho(\xi)$ according to the conditions in Eq. (5) with $\epsilon\neq 0$. The relations in Eq. (5) leads to the next expression for the velocity $v=1/q$ defined in the retarded frame,
$v=\frac{8\epsilon^{2}}{\sigma(\sigma^{2}-4\alpha\epsilon)}.$ (6)
The latter relation shows that the dispersion parameters $\beta_{i}$ $(i=2,3,4)$ included in the parameters $\alpha$, $\sigma$ and $\epsilon$ have a major influence on the wave velocity. This implies that the velocity of the pulse can be controlled by an appropriate manipulation of the dispersion parameters.
With use of the relations in Eq. (5), the system of Eqs. (3) and (4) reduces to the ordinary nonlinear differential equation,
$\epsilon\frac{d^{4}\rho}{d\xi^{4}}+b\frac{d^{2}\rho}{d\xi^{2}}-c\rho+\gamma\rho^{3}=0,~{}~{}~{}~{}~{}~{}~{}$ (7)
where the parameters $b$ and $c$ are
$b=\frac{3\sigma^{2}}{8\epsilon}-\alpha,~{}~{}~{}~{}c=\kappa+\frac{\sigma^{2}}{16\epsilon^{2}}\left(\frac{3\sigma^{2}}{16\epsilon}-\alpha\right).$ (8)
Equation (7) describes the evolution of the field amplitude in the highly dispersive optical fiber. Now the natural question is whether the amplitude equation (7) possesses exact pulse-train solutions without any specified condition on the coefficients or not. In following, we have identified novel families of pulse-train solutions with some interesting properties for this evolutional equation in presence of all physical parameters.
III Results and discussion
In this section, we demonstrate that new kinds of pulse-trains can exist in the fiber system governed by Eq. (1). Remarkably, we find that these propagating waves possess distinct amplitudes and consequently different powers, which is highly desired from a practical point of view.
III.1 Pulse-train solutions
Here we show for the first time that Eq. (7) allows for exact pulse–train solutions which are composed of the product of Jacobi elliptic functions. We find that this type of pulse shape can exist for Eq. (1) in three different classes.
Figure 1: Intensity profiles of pulse-train solutions with parameters $\gamma=2$, $k=0.9$, $\xi_{0}=0$ (a) pulse-train solution (16) when $\alpha=-1,$ $\sigma=-0.25,$ $\epsilon=-0.25$ (b) pulse-train solution (25) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5$ (c) pulse-train solution (34) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5.$
Figure 2: Intensity profiles of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 3: Evolution of pulse-train solutions with parameter $k=0.9$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 4: Evolution of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 5: Dependence of the velocity of pulse-trains $v$ on the group velocity dispersion parameter $\alpha$ for $\sigma=1$ and different values of $\epsilon$, $\epsilon=-0.25$ (thick line), $\epsilon=-0.3$ (dashed line) and $\epsilon=-0.35$ (dotted line).
1. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{dn}\left(x,k\right)$-type
We introduce an appropriate ansatz for the propagating waveform solution of Eq. (7) as,
$\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{dn}\left(w(\xi-\xi_{0}),k\right),$ (9)
where $\xi_{0}$ is an arbitrary constant while $\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)$ and $\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)$ are the cosine and third kind of Jacobian elliptic functions of modulus $k$ taking values $10<k<$ 1.
Substituting the ansatz (9) into Eq. (7) and setting the coefficients of independent terms equal to zero, we obtain the following algebraic equations:
$\epsilon w^{4}(k^{4}-46k^{2}+61)-bw^{2}(k^{2}-5)=c,$ (10)
$60\epsilon w^{4}k^{2}(k^{2}-3)-6bw^{2}k^{2}+\gamma(k^{2}-1)P^{2}=0,$ (11)
$0120\epsilon w^{4}k^{2}+\gamma P^{2}=$ 0 . (12)
By solving the parametric equations (11) and (12), we get
$w=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{5\epsilon^{2}(k^{2}+1)}},~{}~{}~{}~{}P=\pm\frac{k}{k^{2}+1}\sqrt{\frac{-6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (13)
with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Further substitution of Eqs. (13) into Eq. (10) yields the parameter $c$ as
$c=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right)^{2},$ (14)
The wave number $\kappa$ follows from Eqs. (8) and (14) as
$\kappa=\frac{(11k^{4}-86k^{2}+11)}{100\epsilon^{3}(k^{2}+1)^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (15)
Incorporating these results into Eq. (2), we obtain the following periodic wave solution to the extended NLSE (1):
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{dn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (16)
where $\zeta=\tau-v^{-1}z-\xi_{0}$, with $\xi_{0}$ being the position of the wave at $z=0$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
Physically, Eq. (16) describes a periodic wave with amplitude $P$ and inverse temporal width $w$ depending on all order of dispersion as well as nonlinearity.
We now introduce the quantity $\tau_{0}=1/w$ which represents the duration of the pulse [35]. Hence, we find that the relation between the pulse amplitude and duration is given by
$P\tau_{0}^{2}=-2k\left(-\Omega\right)^{1/2}.$ (17)
where the parameter $\Omega=30\epsilon/\gamma$ has been introduced for brevity.
2. Pulse-trains of the $\mathrm{cn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
We now make an ansatz for Eq. (7) as,
$\rho(\xi)=P\,\mathrm{cn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (18)
where $\mathit{\mathrm{sn}}$ is Jacobi elliptic sine function of modulus $k$ which is in the range $10<k<$ 1.
Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
$\epsilon w^{4}(61k^{4}-76k^{2}+16)+bw^{2}(5k^{2}-4)=c,$ (19)
$60\epsilon w^{4}k^{2}(3k^{2}-2)+6bw^{2}k^{2}+\gamma P^{2}=0,$ (20)
$0120\epsilon\eta^{4}k^{2}-\gamma P^{2}=$ 0 . (21)
Thus we have the following relations for the pulse inverse width $w$ and amplitude $\lambda$,
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(4-3k^{2})}},~{}~{}~{}~{}P=\pm\frac{k}{4-3k^{2}}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (22)
with $\gamma\epsilon>0$ and $3\sigma^{2}>8\alpha\epsilon$. Additionally, Eqs. (18) and (21) yield the parameter $c$as
$c=\frac{b^{2}(244k^{2}-89k^{4}-144)}{100\epsilon(4-3k^{2})^{2}},$ (23)
Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
$\kappa=\frac{(244k^{2}-89k^{4}-144)}{100\epsilon^{3}(4-3k^{2})^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (24)
These equations yield the periodic wave solution as
$\psi(z,\tau)=P\,\mathrm{cn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (25)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Noting that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by Eqs. (5) and (6), respectively.
For this solution, we find that the relation between the pulse amplitude and duration takes the form,
$P\tau_{0}^{2}=2k\left(\Omega\right)^{1/2}.$ (26)
3. Pulse-trains of the $\mathrm{dn}\left(x,k\right)\mathrm{sn}\left(x,k\right)$-type
Next we make an ansatz for Eq. (7) as,
$\rho(\xi)=P\,\mathrm{dn}\left(w(\xi-\xi_{0}),k\right)\mathrm{sn}\left(w(\xi-\xi_{0}),k\right),$ (27)
Inserting this ansatz into Eq. (7) and equating the coefficients of independent terms, one obtains
$\epsilon w^{4}(16k^{4}-76k^{2}+61)+bw^{2}(5k^{2}-4)=c,$ (28)
$60\epsilon w^{4}k^{2}(2k^{2}-3)-6bw^{2}k^{2}+\gamma P^{2}=0,$ (29)
$0120\epsilon w^{4}k^{2}-\gamma P^{2}=$ 0 . (30)
Thus we have the following relations for the pulse inverse width $\mu$ and amplitude $R$,
$w=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}(2k^{2}-1)}},~{}~{}~{}~{}P=\pm\frac{k}{2k^{2}-1}\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right).$ (31)
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$ when $11/\sqrt{2}<k<$ 1 and for $8\alpha\epsilon>3\sigma^{2}$ when $0<k<1/\sqrt{2}$.
Moreover, Eqs. (18) and (21) yield the parameter $c$as
$c=\frac{b^{2}(64k^{2}-64k^{4}+11)}{100\epsilon(2k^{2}-1)^{2}},$ (32)
Hence it follows from Eqs. (8) and (22) that the wave number $\kappa$ in the periodic wave solution is given by
$\kappa=\frac{(64k^{2}-64k^{4}+11)}{100\epsilon^{3}(2k^{2}-1)^{2}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (33)
These equations yield the periodic wave solution as
$\psi(z,\tau)=P\,\mathrm{dn}\left(w\zeta,k\right)\mathrm{sn}\left(w\zeta,k\right)\exp[i(\kappa z-\delta\tau+\theta)],$ (34)
where $\xi=\tau-v^{-1}z-\xi_{0}$. Note that the frequency shift $\delta$ and velocity $v$ in this periodic solution are given by the relations (5) and (6), respectively.
For this solution, we find that the relation between the pulse amplitude and duration takes the form,
$P\tau_{0}^{2}=2k\left(\Omega\right)^{1/2}.$ (35)
On comparing Eqs. (16), (25) and (34), we infer that under the influence of second-, third-, and fourth-order dispersions and Kerr nonlinearity the fiber medium allows the propagation of distinct types of pulse-trains which present a relation between the pulse amplitude and duration determined by the sign of the joint parameter $\Omega=30\epsilon/\gamma$ only. To our knowledge, the pulse-trains obtained above for the extended NLSE (1) are firstly reported in this paper.
Figure 1(a) depicts the intensity profile of the pulse-train solutions (16) of Eq. (1) for the parameters values: $\alpha=-1,$ $\gamma=2,$ $\sigma=-0.25,$ and $\epsilon=-0.25.$ The results for the pulse-train solutions (25) and (34) are illustrated in Figs. 1(b) and 1(c) for the values $\alpha=0.49,$ $\gamma=2,$ $\sigma=1,$ and $\epsilon=0.5$. Here the position $\xi_{0}$ of the waveform solutions at $z=0$ is choosen to be equal to zero. Also, the value of the elliptic modulus $k$ is taken as $k=0.9$. In Fig. 2, we show the evolution of these pulse-train solutions in the fiber medium. From these figures, we can see that all these pulse-trains are positive and exhibit an oscillating character. In addition, we can see that the pulse-train solutions (25) and (34)  present two peaks in each periodic unit, markedly different from the pulse train solution (16) which has a one peak structure.
It is also interesting to see from Figs. (3) and (4) that the shape of the pulse-train solution (25) is similar to that for the pulse-train solution (34) for the value of the elliptic modulus $k=0.999$.
The key observation is that the amplitudes, widths and wavenumbers of the derived pulse-train solutions are different whereas their velocity is the same. More importantly, the relation (6) shows that the velocity of the pulse-trains can be significantly reduced for suitable parameters of the waveguiding media, which may find application in developing slow light systems. In Fig. 5, we present the dependence of velocity on the group velocity dispersion parameter $\alpha$ of the fiber for three different values of the quartic dispersion parameter $\epsilon$. It can clearly be seen that the velocity decreases intensely in the magnitude with the increase of dispersion parameter $\alpha$. The decrease of velocity implies that the slowing motion of pulse-trains is possible in the transmitting medium when considering large dispersion parameter $\alpha$. It is also observed that the pulse velocity has large magnitude for small negative values of $\epsilon$. Hence, one concludes that the velocity of pulse-trains can be controlled by judicious choice of the dispersion parameters.
III.2 Quartic and dipole soliton solutions
We now focus on the long-wave limit of the new pulse-trains obtained above corresponding to the modulus $k=1$. In such a case, the periodic solution (16) degenerates to the quartic soliton solution [6]:
$\psi(z,\tau)=P_{0}~{}\mathrm{sech}^{2}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (36)
where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse are given by
$w_{0}=\frac{1}{4}\sqrt{\frac{8\alpha\epsilon-3\sigma^{2}}{10\epsilon^{2}}},$ (37)
$P_{0}=\pm\sqrt{\frac{-3}{10\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (38)
$\kappa=-\frac{4}{25\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (39)
with $\gamma\epsilon<0$ and $8\alpha\epsilon>3\sigma^{2}$. Also the velocity $v$ and frequency shift $\delta$ in this quartic soliton solution are defined by Eqs. (5) and (6).
In addion, when $k\rightarrow 1$, the periodic solution (25) transforms into our dipole soliton solution [7]:
$\psi(z,\tau)=P_{0}~{}\mathrm{sech}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\mathrm{th}\left[w_{0}\left(\tau-v^{-1}z-\xi_{0}\right)\right]\exp[i(\kappa z-\delta\tau+\theta)],$ (40)
where the inverse temporal width $w_{0}$, amplitude $P_{0}$ and wave number $\kappa$ of the soliton pulse take the form
$w_{0}=\frac{1}{4}\sqrt{\frac{3\sigma^{2}-8\alpha\epsilon}{5\epsilon^{2}}},$ (41)
$P_{0}=\pm\sqrt{\frac{6}{5\gamma\epsilon}}\left(\frac{3\sigma^{2}}{8\epsilon}-\alpha\right),$ (42)
$\kappa=\frac{11}{100\epsilon^{3}}\left(\frac{3\sigma^{2}}{8}-\alpha\epsilon\right)^{2}-\frac{\sigma^{2}}{16\epsilon^{3}}\left(\frac{3\sigma^{2}}{16}-\alpha\epsilon\right).$ (43)
with $\gamma\epsilon>0$ and $8\alpha\epsilon<3\sigma^{2}$. The velocity $v$ and frequency shift $\delta$ in this dipole soliton solution are given by Eqs. (5) and (6).
It is interesting to note that if one considers the limiting case with $k=1$, the periodic wave solution (34) also degenerates to the same soliton-like solutions (40).
IV Conclusion
In conclusion, we have discovered three classes of pulse-trains whose amplitudes, widths and wavenumbers are different with the distinctive property of exhibiting a relation between the amplitude and duration determined by the sign of a joint parameter uniquely. We have found that these structures have the same velocity, which depends only on the three dispersion parameters. It is demonstrated that higher-order dispersion may lead to the formation of double-humped pulse-trains besides the ones taking the usual form with single peak in each periodic unit. It is also shown that the pulse-trains are obtained without necessarily assuming a specified condition on the fiber parameters. The results presented here may be useful for experimental realization of pulse-trains in optical fibers and may help in stimulating more research in understanding their optical transmission properties.